# Detecting Adversarial Data via Perturbation Forgery

### v2

@@ Abstract
As a defense strategy against adversarial attacks, adversarial detection aims to identify and filter out adversarial data from the data flow based on discrepancies in distribution and noise patterns between natural and adversarial data. Although previous detection methods achieve high performance in detecting gradient-based adversarial attacks, new attacks based on generative models with imbalanced and anisotropic noise patterns evade detection. Even worse, existing techniques either necessitate access to attack data before deploying a defense or incur a significant time cost for inference, rendering them impractical for defending against newly emerging attacks that are unseen by defenders. In this paper, we explore the proximity relationship between adversarial noise distributions and demonstrate the existence of an open covering for them. By learning to distinguish this open covering from the distribution of natural data, we can develop a detector with strong generalization capabilities against all types of adversarial attacks. Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting unseen gradient-based, generative-based, and physical adversarial attacks, while remaining agnostic to any specific models. Comprehensive experiments conducted on multiple general and facial datasets, with a wide spectrum of attacks, validate the strong generalization of our method.
Table 1: Properties of our method vs Other Adversarial Detection: We summarize four metrics for adversarial detection. “Detect Attacks” means what kind of attacks the method is designed to detect. “Attack-Agnostic” describes whether the detection can detect unseen attack types. “Model-Agnostic” describes whether the method is independent of the protected model. Finally, “Time Cost” measures the time cost of a detector during inference.
@@
Methods Detect Attacks Attack-Agnostic Model-Agnostic Time Cost
LID (Ma et al. 2018) Gradient $\times$ $\times$ $\ast$
LiBRe (Deng et al. 2021) Gradient $\times$ $\ast$
[…]
SPAD (Wang et al. 2023a) Gradient + GAN $\ast$
Ours Gradient + GAN + Diffusion $\ast$
Introduction
@@
Numerous studies have demonstrated the effectiveness of deep neural networks (DNNs) in various tasks (Shi et al. 2020). However, it is also well-known that DNNs are vulnerable to adversarial attacks, which generate adversarial data by adding imperceptible perturbations to natural images, potentially causing the model to make abnormal predictions. This vulnerability reduces the reliability of deep learning systems, highlighting the urgent need for defense techniques. Some approaches employ adversarial training techniques (Zhang et al. 2019), which incorporate adversarial data into the training process to inherently bolster the model’s immunity against adversarial noise (adv-noise). Other methods utilize adversarial purification (Yoon, Hwang, and Lee 2021) to cleanse the input data of adv-noise before model inference. Unfortunately, adversarial training often results in diminished classification accuracy on adversarial data and may sacrifice performance on natural data. While purification techniques may seem more reliable, the denoising process can inadvertently smooth the high-frequency texture of clean images, leading to suboptimal accuracy on both natural and adversarial images.
@@
Another branch of adversarial defense is adversarial detection (Feinman et al. 2017; Ma et al. 2018), which aims to filter out adversarial data before they are fed into target systems. These methods rely on discerning the discrepancies between adversarial and natural distributions, thereby maintaining the integrity of target systems while achieving high performance in defending against adversarial attacks. However, existing adversarial detection approaches primarily train specialized detectors tailored for specific attacks or classifiers (Deng et al. 2021), which may not generalize well to unseen advanced attacks.
@@
Some recent approaches (Zhang et al. 2023; Wang et al. 2023a) have attempted to enhance the generalization of adversarial detectors by leveraging diffusion models and manually designed pseudo-noise. Nevertheless, as shown in Table 1, this improvement often comes at the cost of increased inference time, and even well-trained detectors still struggle with adversarial data produced by generative models such as generative adversarial networks (GANs) and diffusion models. Specifically, generative-based adversarial attacks, such as M3D (Zhao et al. 2023) and Diff-PGD (Xue et al. 2024), are recently developed techniques designed to generate more natural adversarial data. Unlike traditional gradient-based attacks like PGD (Madry et al. 2017), generative-based attacks tend to create imbalanced and anisotropic perturbations that are intense in high-frequency and salient areas but mild in low-frequency and background areas. These perturbations circumvent all current adversarial detection methods, posing a new challenge to the field. Therefore, there is an urgent need for a low-time-cost detection method with strong generalization performance against both unseen gradient-based attacks and unseen generative-based attacks.
@@
In this paper, we investigate the association between the distributions of different adv-noises. As shown in Figure 1, by modeling noise as multivariant Gaussian distributions, we define the proximity relationship of the noise distributions in a metric space and establish a perturbation method to construct the proximal distributions. Based on these assumptions and definitions, we demonstrate that all adv-noise distributions are proximal using the Wasserstein distance, and we deduce a corollary that an open covering of the adv-noise distributions exists. We then propose the core idea of this paper: by learning to distinguish the open covering of adv-noise distributions from the natural distribution, we obtain a detector with strong generalization performance against all types of unseen attacks.
@@
Figure 1: Illustration of the proximity of adv-noise distributions and perturbed noise distributions. Left: all the distributions of adv-noise and the perturbed distributions are in a $\epsilon$-ball centered on a given adv-noise distribution. Right: by continuously perturbing the commonly used adv-noise distribution, we obtain an open covering of distributions of adversarial noise.
@@
Based on this idea, we propose Perturbation Forgery, a pseudo-adversarial data generation method designed to train a detector with strong generalization capabilities for detecting all types of adversarial attacks. As shown in Figure 2, Perturbation Forgery comprises noise distribution perturbation, sparse mask generation, and pseudo-adversarial data (pseudo-adv-data) production. By continuously perturbing the noise distribution of a commonly used attack, such as FGSM (Goodfellow, Shlens, and Szegedy 2014), we obtain a nearly complete family of distributions that form the aforementioned open covering. Simultaneously, sparse masks are generated using attention maps and saliency maps of the natural data, converting the sampled global noises into local forms to help the detector focus on local noise patterns so as to detect generative-based adv-noise. Sampling noises from the perturbed noise distributions, we sparsify these noises with the sparse masks and inject them into natural data. This process converts half of the natural data into pseudo-adv-data, which are then combined with the remaining uncontaminated data to train the binary adversarial detector. To the best of our knowledge, we are the first to empower a detector with the robust ability to detect all types of adversarial attacks, including gradient-based, GAN-based, diffusion-based, and physical attacks, with minor time cost.
@@
$\bullet$ Based on our assumptions and definitions, we deduce the existence of an open covering for adv-noise distributions and propose training a detector to distinguish this open covering from the natural data distribution, thereby achieving strong generalization.
$\bullet$ Building on our deductions, we propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production. This method is designed to train a binary detector with strong generalization performance against all types of unseen attacks, including gradient-based, GAN-based, diffusion-based, and physical adversarial data, with minor time cost.
@@
Related Work
Adversarial Attack
@@
Adversarial attacks (Goodfellow, Shlens, and Szegedy 2014; Madry et al. 2017; Kurakin, Goodfellow, and Bengio 2016) attempt to fool a classifier into changing prediction results by injecting subtle perturbations to the input data while maintaining imperceptibility from human eyes. In an effort to improve transferability under the black-box setting, DIM (Wu et al. 2021), SIM (Lin et al. 2019), and VIM (Wang and He 2021) are recent attacks and achieve a high attack rate against several defense techniques.
@@
Recently, another branch of attack uses generative models to make adversarial data with a more natural style and imbalanced noise compared to gradient-based attacks. Among them, CDA (Naseer et al. 2019), TTP (Naseer et al. 2021), and M3D (Zhao et al. 2023) utilize GAN to craft adversarial data and achieve higher efficiency when attacking large-scale datasets. Diff-Attack (Chen et al. 2023) and Diff-PGD (Xue et al. 2024) employ the diffusion model as a strong prior to better ensure the realism of generated data. Designed for perturbing facial recognition systems, AdvMakeup (Yin et al. 2021) and AMT-GAN (Hu et al. 2022) are presented to generate imperceptible adversarial images of the face.
@@
In this paper, a wide spectrum of gradient-based generative-based and adversarial attacks on ImageNet and facial datasets is used to for evaluation.
Adversarial Detection
@@
As one of the technical solutions to ensure the safety of DNNs, there has been a significant amount of exploration in the field of adversarial detection (Ma et al. 2018). LiBRe (Deng et al. 2021) leverages Bayesian neural networks to endow pre-trained DNNs with the ability to defend against unseen adversarial attacks, SPAD (Wang et al. 2023a) leverages a data augmentation method to detect adversarial face images from unseen attack, while EPSAD (Zhang et al. 2023) incorporates a diffusion model to detect adversarial data with minimal perturbation magnitude, albeit at the cost of inference time.
@@
While current methods have achieved remarkable improvements in adversarial detection, generative-based adversarial attacks can easily circumvent them, posing a new challenge. To address the problems above, our method is proposed to detect all types of unseen attacks, including gradient-based, generative-based, and physical attacks.
Proximity of Adversarial Noise Distribution
@@
In this section, we provide an overview of preliminary concepts related to adversarial data generation and explore the relationship between adversarial noise distributions. We define the proximity of noise distributions and examine the imperceptibility of adversarial noise. Following this, we present the core idea of this paper: by learning to distinguish the open covering of adversarial noise distributions from the natural distribution, we develop a detector with strong generalization capabilities against all types of unseen attacks.
Adversarial data generation
@@
Given a DNN $f$ that works on dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}^{n}_{i=1}$ with $(\mathbf{x}_{i}\in\mathcal{X},y_{i}\in\mathcal{C})$ being (sample, ground-truth) pair, an adversarial example $\hat{\mathbf{x}}$ regarding sample $\mathbf{x}$ with perturbation magnitude $\epsilon$ is calculated as:
$\hat{\mathbf{x}}=\mathop{\mathrm{argmax}}_{\hat{\mathbf{x}}\in\mathbb{B}(\mathbf{x},\epsilon)}\mathcal{L}(f(\hat{\mathbf{x}}),y),$ (1)
@@
where $\mathcal{L}$ is some loss function and $\mathcal{B}(\mathbf{x},\epsilon)=\{\mathbf{x^{\prime}}|\ell^{p}(\mathbf{x},\mathbf{x^{\prime}})<\epsilon\}$ restricts the perturbation under $\ell^{p}$ norm. Following (Goodfellow, Shlens, and Szegedy 2014), we simply denote adversarial sample as $\hat{\mathbf{x}}=\mathbf{x}+\boldsymbol{\eta}$.
Proximity of noise distribution
@@
Due to the subtle and imperceptible nature of adv-noise, adversarial data and natural data appear visually similar. Thus, we assume that this similarity also exists between the noises of different adversarial attacks. To investigate the similarity of these noises, we provide the following definition, which depicts a close relationship between two independent noise distributions:
@@
Let $P$ and $Q$ be the distribution functions of two independent noise samples $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. We call them proximal noise distributions under $\varepsilon$ if we have a metric $d(\cdot,\cdot)$ and $\exists\ \varepsilon>0$, s.t.
$d(P,Q)<\varepsilon.$ (2)
A common assumption for instance feature distribution is to model it as a Gaussian distribution (Du et al. 2022). Obtained by subtracting from clean data and adversarial data, adv-noises naturally form a Gaussian distribution. Therefore, we assume that noise r.v $\boldsymbol{\eta}$ satisfied truncated Gaussian distribution $\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ with it’s truncated interval is $[-\epsilon,\epsilon]^{d}$. Built upon Definition 1, we derive the following theorem about the association between the distributions of different adv-noises:
@@
Theorem 1 tells us: 1) All adversarial noise distributions are proximal noise distributions. 2) If the parameters of two noise distributions are sufficiently close to those of a given distribution, we can regard these two distributions as proximal noise distributions.
Figure 2: Illustration of Perturbation Forgery. Before training, we estimate the noise distribution of a commonly used attack and continuously perturb it in each batch. After that, we sample noises from the perturbed distribution and convert them into local noise using sparse masks. Finally, pseudo-adversarial data is generated by adding these local noises to the natural samples.
@@
We can further deduce several corollaries of the theorem. 1) All noise distributions proximal to the known adversarial distribution $p(\boldsymbol{\eta})$ form an open $\varepsilon$-ball centered on $p(\boldsymbol{\eta})$, denoting as $\mathcal{P}_{\varepsilon}=\{p(\boldsymbol{\eta}_{i})|d(p(\boldsymbol{\eta}_{i}),p(\boldsymbol{\eta}))<\varepsilon)\}$, and also form a metric space $(\mathcal{P}_{\varepsilon},d)$. 2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$. 3) Based on the properties of the metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists an open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in J}$.
@@
Based on the corollaries, we deduce the core idea of this paper: Let $p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})=\mathcal{N}(\mu+\epsilon_{\mu i},(\sigma+\epsilon_{\sigma i})^{2}\mathbf{I}_{d}),i=1,2,\dots,n$ denote generated proximal noise distributions of the given adv-noise distribution $p(\boldsymbol{\eta})=\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ where $\epsilon_{\mu i},\epsilon_{\sigma i}$ are randomly sampled from uniform distributions and restricted by $\epsilon_{\mu i},\epsilon_{\sigma i}<\varepsilon$. As shown in Figure 1, when $n\rightarrow\infty$, we can always obtain a distribution set $\mathcal{P}_{c}=\{p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})\}_{i=1}^{\infty}$ that forms an open covering of the adversarial distribution set $\mathcal{P}_{a}$. Trained with the open covering, the detector can distinguish the distributions of natural data and adversarial data, thereby having the functionality to detect unseen adversarial attacks.
@@
For proofs of theorem and corollaries, please refer to the supplementary materials.
Perturbation Forgery

### v4

# Detecting Adversarial Data Using Perturbation Forgery

@@ Abstract
As a defense strategy against adversarial attacks, adversarial detection aims to identify and filter out adversarial data from the data flow based on discrepancies in distribution and noise patterns between natural and adversarial data. Although previous detection methods achieve high performance in detecting gradient-based adversarial attacks, new attacks based on generative models with imbalanced and anisotropic noise patterns evade detection. Even worse, the significant inference time overhead and limited performance against unseen attacks make existing techniques impractical for real-world use. In this paper, we explore the proximity relationship among adversarial noise distributions and demonstrate the existence of an open covering for these distributions. By training on the open covering of adversarial noise distributions, a detector with strong generalization performance against various types of unseen attacks can be developed. Based on this insight, we heuristically propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production, to train an adversarial detector capable of detecting any unseen gradient-based, generative-based, and physical adversarial attacks. Comprehensive experiments conducted on multiple general and facial datasets, with a wide spectrum of attacks, validate the strong generalization of our method. [^1] Code at https://github.com/cc13qq/PFD
Table 1: Properties of our method vs Other Adversarial Detection: We summarize four metrics for adversarial detection. “Detect Attacks” specifies the types of attacks each method is designed to detect. “Model-Agnostic” indicates whether the detection method operates independently of the protected model. “Unseen-Attack Detection” assesses the method’s generalization to detect previously unseen attacks, with the reported value representing the AUROC score for detecting the MIFGSM attack on the ImageNet dataset. Lastly, “Time Overhead” quantifies the time required for the detector to process 100 samples from the ImageNet dataset.
@@
Methods Detect Attacks Model-Agnostic Unseen-Attack Detection Time Overhead
LID [28] Gradient $\times$ 0.9146 1.80s
LiBRe [8] Gradient $\times$ 0.8725 2.56s
SPAD [40] Gradient + GAN 0.9820 4.56s
EPSAD [49] Gradient 0.9918 396.81s
Ours Gradient + GAN + Diffusion 0.9931 4.85
1 Introduction
@@
Numerous studies have demonstrated the effectiveness of deep neural networks (DNNs) in various tasks [36]. However, it is also well-known that DNNs are vulnerable to adversarial attacks, which generate adversarial data by adding imperceptible perturbations to natural images, potentially causing the model to make abnormal predictions. This vulnerability reduces the reliability of deep learning systems, highlighting the urgent need for defense techniques. Some approaches employ adversarial training techniques [48], which incorporate adversarial data into the training process to inherently bolster the model’s immunity against adversarial noise (adv-noise). Other methods utilize adversarial purification [47] to cleanse the input data of adv-noise before model inference. Unfortunately, adversarial training often results in diminished classification accuracy on adversarial data and may sacrifice performance on natural data [1]. While purification techniques may seem more reliable, the denoising process can inadvertently smooth the high-frequency texture of clean images, leading to suboptimal accuracy on both natural and adversarial images.
@@
Another branch of adversarial defense is adversarial detection [12, 28], which aims to filter out adversarial data before they are fed into target systems. These methods rely on discerning the discrepancies between adversarial and natural distributions, thereby maintaining the integrity of target systems while achieving high performance in defending against adversarial attacks. However, existing adversarial detection approaches primarily train specialized detectors tailored for specific attacks or classifiers [8], which may not generalize well to unseen advanced attacks [39, 2].
@@
Some recent approaches [49, 40] have attempted to enhance the generalization of adversarial detectors by leveraging diffusion models and manually designed pseudo-noise. Nevertheless, as shown in Table 1, this improvement often comes at the expense of increased inference overhead, and even well-trained detectors still struggle with adversarial data produced by generative models such as generative adversarial networks (GANs) and diffusion models. Specifically, generative-based adversarial attacks, such as M3D [51] and Diff-PGD [44], are recently developed techniques designed to generate more natural adversarial data. Unlike traditional gradient-based attacks like PGD [29], generative-based attacks tend to create imbalanced and anisotropic perturbations that are intense in high-frequency and salient areas but mild in low-frequency and background areas. These perturbations make them harder to detect with current methods, posing a new challenge to the field of adversarial detection. Therefore, there is an urgent need for a low-time-cost detection method with strong generalization performance against both unseen gradient-based attacks and unseen generative-based attacks.
@@
In this paper, we investigate the association between the distributions of different adv-noises. As shown in Figure 1, by modeling noise as multivariate Gaussian distributions, we define the proximity relationship of the noise distributions in a metric space and establish a perturbation method to construct the proximal distributions. Based on the assumption and definition, we demonstrate that all adv-noise distributions are proximal using the Wasserstein distance, and we deduce a corollary that an open covering of the adv-noise distributions exists. We then propose the core idea of this paper: by training on the open covering, we obtain a detector with strong generalization performance against various types of unseen attacks.
@@
Figure 1: Illustration of the proximity of adversarial noise distributions and perturbed noise distributions. Left: all the distributions of adversarial noise and the perturbed distributions are in a $\varepsilon$-ball centered on a given adversarial noise distribution. Right: by continuously perturbing the given adversarial noise distribution, we obtain an open covering of distributions of adversarial noise.
@@
Based on this idea, we propose Perturbation Forgery, a pseudo-adversarial data (pseudo-adv-data) generation method designed to train a detector with strong generalization capabilities for detecting various types of adversarial attacks. As shown in Figure 2, Perturbation Forgery comprises noise distribution perturbation, sparse mask generation, and pseudo-adversarial data (pseudo-adv-data) production. By continuously perturbing the noise distribution of a commonly used attack, such as FGSM [15], we obtain a nearly complete family of distributions that form the aforementioned open covering. Simultaneously, sparse masks are generated using attention maps and saliency maps of the natural data, converting the sampled global noises into local forms to help the detector focus on local noise patterns so as to detect generative-based adv-noise. By sampling noise from the perturbed noise distributions, we sparsify it using sparse masks and inject it into natural data. This process converts half of the natural data into pseudo-adv-data, replacing actual adversarial data. The pseudo-adv-data is then combined with the remaining uncontaminated data to train the binary adversarial detector.
To the best of our knowledge, we are the first to empower a detector with the robust ability to detect various types of adversarial attacks, including gradient-based, GAN-based, diffusion-based, and physical attacks, with minor inference time overhead.
@@
$\bullet$ Based on our assumption and definition, we deduce the existence of an open covering for adversarial noise distributions and propose training a detector to distinguish this open covering from the natural data distribution, thereby achieving strong generalization in detection.
$\bullet$ Building on our deductions, we propose Perturbation Forgery, which includes noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production. This method is designed to train a binary detector with strong generalization performance against various types of unseen attacks, including gradient-based, GAN-based, diffusion-based, and physical adversarial data, with minor inference time overhead.
@@
2 Related Work
2.1 Adversarial Attack
@@
Adversarial attacks [15] attempt to fool a classifier into changing prediction results by injecting subtle perturbations to the input data while maintaining imperceptibility from human eyes. By introducing iterative perturbation, PGD [29] and BIM [24] generate adversarial data with significantly improved attack performance, while DIM [43], NIM [26], and MIM [9] improve the transferability under the black-box setting. Besides, adaptive attacks [6, 13, 18] are proposed to exploit the weak points of each defense, achieving a high attack rate against several defense techniques.
@@
Recently, another branch of attacks uses generative models to produce adversarial data with a more natural style and imbalanced noise compared to gradient-based attacks. Among them, CDA [31], TTP [30], and M3D [51] utilize GAN to craft adversarial data and achieve higher efficiency when attacking large-scale datasets. Diff-Attack [4] and Diff-PGD [44] employ the diffusion model as a strong prior to better ensure the realism of generated data. Designed for perturbing facial recognition systems, AdvMakeup [46] and AMT-GAN [19] are presented to generate imperceptible adversarial images of the face.
@@
In this paper, a wide spectrum of gradient-based, generative-based, and physical adversarial attacks on ImageNet and facial datasets is used to evaluate the detection performance of our method.
2.2 Adversarial Detection
@@
As one of the technical solutions to ensure the safety of DNNs, there has been a significant amount of exploration in the field of adversarial detection [28]. LiBRe [8] leverages Bayesian neural networks to endow pre-trained DNNs with the ability to defend against unseen adversarial attacks, SPAD [40] leverages a data augmentation method to detect adversarial face images from unseen attack, while EPSAD [49] incorporates a diffusion model to detect adversarial data with minimal attack intensity, albeit at the cost of inference time.
@@
While current methods have achieved remarkable improvements in adversarial detection, generative-based adversarial attacks are harder to detect, posing a new challenge. To address the problems above, our method is proposed to detect various types of unseen attacks, including gradient-based, generative-based, and physical attacks.
3 Proximity of Adversarial Noise Distribution
@@
In this section, we provide an overview of preliminary concepts related to adversarial data generation and explore the relationship between adversarial noise distributions. We define the proximity of noise distributions and present the core idea of this paper: by training on the open covering of adversarial noise distributions, we obtain a detector with strong generalization performance against various types of unseen attacks.
3.1 Adversarial data generation
@@
Given a DNN $f$ that works on dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}^{n}_{i=1}$ with $(\mathbf{x}_{i}\in\mathcal{X},y_{i}\in\mathcal{C})$ being (sample, ground-truth) pair, an adversarial example $\hat{\mathbf{x}}$ regarding sample $\mathbf{x}$ with attack intensity $\epsilon$ is calculated as:
$\hat{\mathbf{x}}=\mathop{\mathrm{argmax}}_{\hat{\mathbf{x}}\in\mathcal{B}(\mathbf{x},\epsilon)}\mathcal{L}(f(\hat{\mathbf{x}}),y),$ (1)
@@
where $\mathcal{L}$ is some loss function and $\mathcal{B}(\mathbf{x},\epsilon)=\{\mathbf{x^{\prime}}|\ell^{p}(\mathbf{x},\mathbf{x^{\prime}})<\epsilon\}$ restricts the perturbation under $\ell^{p}$ norm. Following [15], we simply denote adversarial sample as $\hat{\mathbf{x}}=\mathbf{x}+\boldsymbol{\eta}$.
Figure 2: Illustration of Perturbation Forgery. Before training, we estimate the noise distribution from a commonly used attack, then continuously perturb it in each batch to create an open covering of the adversarial noise distributions. Next, noises sampled from these perturbed distributions are converted into localized noise by applying sparse masks. Finally, pseudo-adversarial data is generated by adding these localized noises to natural samples.
3.2 Proximity of noise distribution
@@
Due to the subtle and imperceptible nature of adv-noise, adversarial data and natural data appear visually similar. Thus, we assume that this similarity also exists between the noises of different adversarial attacks. To investigate the similarity of these noises, we provide the following definition, which depicts a close relationship between two noise distributions.
@@
Let $P$ and $Q$ be the distribution functions of two independent noise samples $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. We call them proximal noise distributions under $\varepsilon$ if we have a metric $d(\cdot,\cdot)$ and $\exists\ \varepsilon>0$, s.t. $d(P,Q)<\varepsilon$.
A common assumption for instance feature distribution is to model it as a Gaussian distribution [10]. Obtained by subtracting from natural data and adversarial data, adv-noises naturally form a Gaussian distribution. Therefore, we assume that noise r.v $\boldsymbol{\eta}\in\mathbb{R}^{d}$ satisfied truncated Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\sigma^{2}\mathbf{I}_{d})$ with it’s truncated interval is $[-\epsilon,\epsilon]^{d}$. Built upon Definition 1, we derive the following theorem about the association between the distributions of different adv-noises:
@@
Theorem 1 tells us: 1) All the adversarial noise distributions under the same $\ell^{p}$ norm are proximal noise distributions. 2) If the parameters of two noise distributions are sufficiently close to those of a given distribution, we can regard these two distributions as proximal noise distributions.
@@
We can further deduce several corollaries of the theorem. 1) All noise distributions proximal to the known adversarial distribution $p(\boldsymbol{\eta})$ form an open $\varepsilon$-ball centered on $p(\boldsymbol{\eta})$, denoting as $\mathcal{P}_{\varepsilon}=\{p(\boldsymbol{\eta}_{i})|d(p(\boldsymbol{\eta}_{i}),p(\boldsymbol{\eta}))<\varepsilon)\}$, and also form a metric space $(\mathcal{P}_{\varepsilon},d)$. 2) The adv-noise distribution set is also located in this ball: $\mathcal{P}_{a}\subset\mathcal{P}_{\varepsilon}$. 3) Based on the properties of the metric space $(\mathcal{P}_{\varepsilon},d)$, for all subset $\mathcal{P}_{i}\subset\mathcal{P}_{\varepsilon}$, there exists an open covering of the subset: $\mathcal{P}_{i}\subset\{G_{c}\}_{c\in\mathbb{I}}$.
@@
Based on the corollaries, we deduce the core idea of this paper: Let $p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})=\mathcal{N}(\boldsymbol{\mu}+\epsilon_{\mu i},(\sigma+\epsilon_{\sigma i})^{2}\mathbf{I}_{d}),i=1,2,\dots,n$ denote generated proximal noise distributions of the given adv-noise distribution $p(\boldsymbol{\eta})=\mathcal{N}(\boldsymbol{\mu},\sigma^{2}\mathbf{I}_{d})$, where $\epsilon_{\mu i},\epsilon_{\sigma i}$ are randomly sampled from uniform distributions and restricted by $\epsilon_{\mu i},\epsilon_{\sigma i}<\varepsilon$. As shown in Figure 1, when $n\rightarrow\infty$, we can always obtain a distribution set $\mathcal{P}_{c}=\{p(\hat{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})\}_{i=1}^{\infty}$ that forms an open covering of the adversarial distribution set $\mathcal{P}_{a}$. By training on the open covering, the detector learns to distinguish between the distributions of natural and adversarial data, enabling it to effectively detect unseen adversarial attacks.
@@
Analysis of distribution assumption, proofs of theorem, and corollaries are provided in the supplementary materials.
4 Perturbation Forgery